# ArchRAG: Attributed Community-based Hierarchical Retrieval-Augmented Generation

### v1

@@ 4.3. Detailed Analysis
In addition, we evaluate the performance of ArchRAG under different LLM backbones and various top-k retrieval content to address question RQ4.
$\bullet$ Effect of LLM backbones. As strong LLMs with hundreds of billions of parameters (e.g., GPT-3.5-turbo) possess enhanced capabilities, our proposed ArchRAG may also benefit from performance improvement. As shown in Table 5, the results of Llama 3.1-8B are similar to those of GPT-3.5-turbo, as Llama3.1’s capabilities are comparable to those of GPT-3.5-turbo (Dubey et al., 2024). Besides, GPT-4o-mini performs better than other LLM backbones because of its exceptional reasoning capabilities.
Table 5. Comparing ArchRAG with other RAG methods on the specific QA tasks under different LLM backbone models.
LLM backbone Multihop-RAG HotpotQA
(Accuracy) (Recall) (Accuracy) (Recall)
Llama3.1-8B 68.8 37.2 65.4 69.2
GPT-3.5-turbo 67.2 31.5 62.8 65.0
GPT-4o-mini 77.3 33.8 69.9 73.8
LABEL:topk_leg
5. Related Work
$\bullet$ RAG approaches. RAG has been proven to excel in many tasks, including open-ended question answering (Jeong et al., 2024; Siriwardhana et al., 2023), programming context (Chen et al., 2024a, 2023), SQL rewrite (Li et al., 2025; Sun et al., 2024), and data cleaning (Naeem et al., 2024; Narayan et al., 2022; Qian et al., 2024). The naive RAG technique relies on retrieving query-relevant contexts from external knowledge bases to mitigate the “hallucination” of LLMs. Recently, many RAG approaches (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024; Sarthi et al., 2024; Wang et al., 2024a; He et al., 2024; Ma et al., 2024, 2024; Li et al., 2024) have adopted graph structures to organize the information and relationships within documents, achieving improved overall retrieval performance. For more details, please refer to the recent survey of graph-based RAG methods (Peng et al., 2024).
$\bullet$ LLM for graph mining. Recent advances in LLMs have offered opportunities to leverage LLMs in graph mining. These include using LLMs for KG construction (Zhu et al., 2024), and employing KG to enhance the LLM reasoning (Wang et al., 2023a, 2024a; Luo et al., 2023; Sun et al., 2023; Mavromatis and Karypis, 2024; Jiang et al., 2023). For instance, KD-CoT (Wang et al., 2023a) retrieves facts from an external knowledge graph to guide the CoT performed by LLMs. RoG (Luo et al., 2023) proposes a planning-retrieval-reasoning framework that retrieves reasoning paths from KGs to guide LLMs conducting faithful reasoning. To capture graph structure, GNNRAG (Mavromatis and Karypis, 2024) adopts a lightweight graph neural network to retrieve elements from KGs effectively. StructGPT (Jiang et al., 2023) and ToG (Sun et al., 2023) treat LLMs as agents that interact with KGs to find reasoning paths leading to the correct answers. Besides, LLM is used to generate code to solve complex graph mining problems, e.g., the shortest path computation (Chen et al., 2024b; Tang et al., 2024; Cao et al., 2024; Zhang et al., 2024a).
$\bullet$ LLM for graph learning. Recently, LLMs have demonstrated remarkable zero-shot capabilities in processing text-attributed graphs. Following this trend, graph foundation models have emerged as a promising direction (Liu et al., 2024a; Zhang et al., 2024b). For instance, Liu et al. (Liu et al., 2024a) proposed a unified model capable of handling various graph ML tasks, such as node classification and link prediction, thereby eliminating the need for task-specific architectures.
6. Conclusion
@@
Appendix A Method details
A.1. LLM-based hierarchical clustering
@@
A.2. More details of C-HNSW
@@
Definition 1 (Hierarchical Navigable Small World (HNSW) (Malkov and Yashunin, 2018)).
@@
$\bullet$ The construction of C-HNSW. The construction of C-HNSW is illustrated in Algorithm 3. The construction algorithm of C-HNSW follows a top-down approach. Using the query process of C-HNSW, we obtain the $M$ nearest neighbors of each node in its layer, and the inter-layer links are continuously updated during this process. When inserting a node $x$ at layer $i$ , if the nearest neighbor at the layer $i+1$ is $c_{j}$, the inter-layer link of $c_{j}$ is updated in the following two cases:
@@ 7                   $c\leftarrow$ get the nearest node from R;
8                   $s\leftarrow$ node in layer $l-1$ via $c$’s inter-layer link;
@@ 15      $\mathcal{H}\leftarrow\mathcal{H}\cup{G_{l}}=(V_{l},E_{l})$
A.3. Complexity analysis of ArchRAG
@@
Given a large text corpus or a large set of text documents with a total of $D$ tokens, the time complexity of the offline indexing process of ArchRAG is $O(I\frac{D}{w}+\frac{1-a^{L}}{1-a}(n*t+I\frac{D}{w}+\pi(m)+n\log n))$, where ${\bf I}$ is the generation time of the LLM for a single inference, $w$ is the specified token size of one chunk, $n$ and $m$ are the number of entities and relations in the extracted KG, $L$ is the height of the resulting hierarchical community structure, and $a$ is the average ratio of the number of nodes between two consecutive layers, with $10<a<$ 1. For a given embedding model, the computation time for the embedding of an entity description is denoted by $t$, while a specific clustering method is typically a function of $m$, represented as $\pi(m)$.
@@
For the corpus $D$, we use the LLM to infer and extract the KG from each chunk of size $w$, resulting in a cost of $O(I\frac{D}{w})$ for constructing the KG. Since the size of all community summaries in a single layer would not exceed the length of the corpus, their LLM inference time is also less than $O(I\frac{D}{w})$. For each layer of clustering, the embedding of each point must be computed, and clustering is performed with time complexity of $\pi(m)$. In the C-HNSW construction, each point requires $O(nlogn)$ time to perform the k-nearest neighbor search and establish connections, which is similar to the proof in (Malkov and Yashunin, 2018). For an L-layer multi-layer graph structure, where the number of nodes decreases by a factor of $a$ between two consecutive layers, the increase in clustering and C-HNSW construction time is given by: $\frac{1-a^{L}}{1-a}$. ∎
@@
Given a large text corpus or a large set of text documents with a total of $D$ tokens, the number of tokens used in the offline indexing process of ArchRAG is $O(D+D\frac{1-a^{L}}{1-a})$, where $L$ is the height of the resulting hierarchical community structure and $a$ is the average ratio of the number of nodes between two consecutive layers, with $10<a<$ 1.
@@
ArchRAG first computes the embedding of the query, which takes $O(e)$ time. For each layer, querying one nearest neighbor takes no more than $O(\log(n))$ time, similar to the proof in (Malkov and Yashunin, 2018). In the Adaptive filtering-based generation, the content of each query is analyzed and inferred, requiring $O(LkI)$ time. Therefore, the total time for the online retrieval query is $O(e+LkI+Lk\log(n))$. ∎
@@
Given a C-HNSW with $L$ layers constructed from a large text corpus or a large set of text documents, the number of tokens used for a single online retrieval query in ArchRAG is $O(kL(c+P)))$, where $k$ is the number of neighbors retrieved at each layer, $c$ is the average length of the retrieved content, and $P$ is the length of the prompt.
@@ B.1. Metrics
$\bullet$ Metrics for specific QA tasks. We choose accuracy as the evaluation metric based on whether the gold answers are included in the model’s generations rather than strictly requiring an exact match, following (Schick et al., 2024; Mallen et al., 2022; Asai et al., 2023). This is because LLM outputs are typically uncontrollable, making it difficult for them to match the exact wording of standard answers. Similarly, we choose recall as the metric instead of precision, as it better reflects the accuracy of the generated responses. Additionally, when calculating recall, we adopt the same approach as previous methods (Gutiérrez et al., 2024; Asai et al., 2023): if the golden answer or the generated output contains “yes” or “no”, the recall for that question is set to 0. Therefore, the recall metric is not perfectly correlated with accuracy.
@@
$\bullet$ Metrics for abstract QA tasks. Following existing works, we use an LLM to generate abstract questions, with the prompts shown in Figure 8, defining ground truth for abstract questions, particularly those involving complex high-level semantics, poses significant challenges. We build on existing works (Edge et al., 2024; Guo et al., 2024) to address this and adopt an LLM-based multi-dimensional comparison method (including comprehensiveness, diversity, empowerment, and overall). We employ a robust LLM, specifically GPT-4o-mini, to rank each baseline against our method. Figure 9 shows the evaluation prompt we use.
@@
Prompt for generating abstract questions Prompt: Given the following description of a dataset: {description} Please identify 5 potential users who would engage with this dataset. For each user, list 5 tasks they would perform with this dataset. Then, for each (user, task) combination, generate 5 questions that require a high-level understanding of the entire dataset. Output the results in the following structure: - User 1: [user description]: : - Task 1: [task description]: : : - Question 1: : - Question 2: : - Question 3: : - Question 4: : - Question 5: - Task 2: [task description]: : - Task 5: [task description]: - User 2: [user description]: : - User 5: [user description]: : Note that there are 5 users and 5 tasks for each user, resulting in 25 tasks in total. Each task should have 5 questions, resulting in 125 questions in total. The Output should present the whole tasks and questions for each user. Output:
Figure 8. The prompt for generating abstract questions.
@@
        "Explanation": "[Briefly summarize why this answer is the overall winner]"
@@
Figure 9. The prompt for the evaluation of abstract QA.
  1. (1)
    Calinski-Harabasz Index (CHI) (Caliński and Harabasz, 1974): A higher value of CHI indicates better clustering results because it means that the data points are more spread out between clusters than they are within clusters. It is an internal evaluation metric where the assessment of the clustering quality is based solely on the dataset and the clustering results and not on external ground-truth labels. The CHI is calculated by between-cluster separation and within-cluster dispersion:
    (5) $CHI=\frac{N-C}{C-1}\frac{\sum^{C}_{i=1}n_{i}||c_{i}-c||^{2}}{\sum^{C}_{i=1}\sum_{x\in C_{i}}||x-c_{i}||^{2}}.$
    $N$ is the number of nodes. $C$ is the number of clusters.$n_{i}$ is the number of nodes in cluster $i$. $C_{i}$ is the $i-th$ cluster. $c_{i}$ is the centroid of cluster $C_{i}$. $c$ is the overall centroid of the datasets. $x$ is the feature of the target node.
  2. (2)
    Cosine Similarity (Sim) (Charikar, 2002): Cosine similarity is a measure of similarity between two non-zero vectors defined in an inner product space. In this paper, for each cluster, we calculate the similarity between the centroid of this cluster and each node in this cluster:
    (6) $Sim=\frac{1}{N}\sum^{C}_{i=1}\sum_{x\in C_{i}}Cosine(x,c_{i}).$
    $N$ is the number of nodes. $C$ is the number of clusters. $C_{i}$ is the $i$-th cluster. $c_{i}$ is the centroid of cluster $C_{i}$. $x$ is the feature of the target node.
    (7) $Cosine(x,y)=\frac{\sum_{i=1}^{n}x_{i}y_{i}}{\sqrt{\sum_{i=1}^{n}x_{i}^{2}}\sqrt{\sum_{i=1}^{n}y_{i}^{2}}}.$
B.2. Implementation details
We implement our ArchRAG in Python, while C-HNSW is implemented in C++ and provides a Python interface for integration. We implement C-HNSW using the FAISS framework and employ the inner product metric to measure the proximity between two vectors. All methods utilize 10 concurrent LLM calls, and to maintain consistency, other parallel computations in the method, such as embedding calculations, also use 10 concurrent threads. All the experiments were conducted on a Linux operating system running on a machine with an Intel Xeon 2.0GHz CPU, 1024GB of memory, and 8 NVIDIA GeForce RTX A5000 GPUs, each with 24 GB of memory. Figure 10 demonstrates the prompt used in Adaptive filtering-based generation. Please refer to our repository (anonymous.4open.science/r/H-CAR-AG-1E0B) to view the detailed prompts.
Prompt for Adaptive filtering-based generation Filger Prompt: # Role You are a helpful assistant responding to questions about data in the tables provided. # Goal Generate a response consisting of a list of key points that respond to the user’s question, summarizing all relevant information in the input data tables. You should use the data provided in the data tables below as the primary context for generating the response. If you don’t know the answer or if the input data tables do not contain sufficient information to provide an answer, just say so. Do not make anything up. Each key point in the response should have the following element: Description: A comprehensive description of the point. Importance Score: An integer score between 0-100 that indicates how important the point is in answering the user’s question. An ’I don’t know’ type of response should have a score of 0. The response should be JSON formatted as follows:

### v3

@@ 21          $Q\leftarrow Q\cup\{x\}$, $K\leftarrow K\cup\{x\}$;
$\bullet$ Impact of Attributed Communities in RAG. We propose a new variant, GraphRAG+AC, which replaces the original Leiden-based communities in GraphRAG with our ACs, while preserving the original Global Search pipeline. As shown in Table 4, this variant results in a significant performance improvement compared to the original approach. Specifically, on the HotpotQA dataset, GraphRAG+AC improves accuracy by 51% compared to GGraphRAG.
Table 4: Results of GraphRAG variants using our ACs.
Method Multihop-RAG HotpotQA
(Acc) (Rec) (Acc) (Rec)
GGraphRAG 45.9 28.4 33.5 42.6
GraphRAG+AC 49.3 31.4 50.6 (51.0% $\uparrow$) 52.8
ArchRAG 68.8 37.2 65.4 69.2
We further test ArchRAG under different LLM backbones, various top-$k$ retrieval settings, and evaluate the community quality of its LLM-based hierarchical clustering. ArchRAG consistently achieves state-of-the-art performance across all backbones, remains robust under varying top-$k$ values, and demonstrates high-quality community detection. We also provide additional case studies of our ArchRAG, which illustrate the detailed workflow. Detailed experiments are provided in the appendix.
Conclusion
@@
Appendix A Method details of ArchRAG
LLM-based hierarchical clustering
@@ Appendix A Method details of ArchRAG
More details of C-HNSW
@@
Definition 1 (Hierarchical Navigable Small World (HNSW) (Malkov and Yashunin 2018)).
@@
$\bullet$ The construction of C-HNSW. A naive approach to build the C-HNSW index is to build nodes first and then establish the two types of links by finding the nearest neighbors of each node. However, the process of finding the nearest neighbor is costly. To accelerate the construction, we propose a top-down approach by borrowing the idea of HNSW construction. The construction of C-HNSW is illustrated in Algorithm 3. The construction algorithm of C-HNSW follows a top-down approach. Using the query process of C-HNSW, we obtain the $M$ nearest neighbors of each node in its layer, and the inter-layer links are continuously updated during this process. When inserting a node $x$ at layer $i$ , if the nearest neighbor at the layer $i+1$ is $c_{j}$, the inter-layer link of $c_{j}$ is updated in the following two cases:
@@ 7       $c\leftarrow$ get the nearest node from R;
8       $s\leftarrow$ node in layer l1l-1 via $c$’s inter-layer link;
@@ 15 $\mathcal{H}\leftarrow\mathcal{H}\cup{G_{l}}=(V_{l},E_{l})$
Complexity analysis of ArchRAG
@@
Given a large text corpus or a large set of text documents with a total of $D$ tokens, the time complexity of the offline indexing process of ArchRAG is $O(I\frac{D}{w}+\frac{1-a^{L}}{1-a}(n*t+I\frac{D}{w}+\pi(m)+n\log n))$, where $\bf I$} is the generation time of the LLM for a single inference, $w$ is the specified token size of one chunk, $n$ and $m$ are the number of entities and relations in the extracted KG, $L$ is the height of the resulting hierarchical community structure, and $a$ is the average ratio of the number of nodes between two consecutive layers, with $0<a<1$. For a given embedding model, the computation time for the embedding of an entity description is denoted by $t$, while a specific clustering method is typically a function of $m$, represented as $\pi(m)$.
@@
For the corpus $D$, we use the LLM to infer and extract the KG from each chunk of size $w$, resulting in a cost of $O(I\frac{D}{w})$ for constructing the KG. Since the size of all community summaries in a single layer would not exceed the length of the corpus, their LLM inference time is also less than $O(I\frac{D}{w})$. For each layer of clustering, the embedding of each point must be computed, and clustering is performed with time complexity of $\pi(m)$. In the C-HNSW construction, each point requires $O(nlogn)$ time to perform the k-nearest neighbor search and establish connections, which is similar to the proof in (Malkov and Yashunin 2018). For an L-layer multi-layer graph structure, where the number of nodes decreases by a factor of $a$ between two consecutive layers, the increase in clustering and C-HNSW construction time is given by: $\frac{1-a^{L}}{1-a}$. ∎
@@
Given a large text corpus or a large set of text documents with a total of $D$ tokens, the number of tokens used in the offline indexing process of ArchRAG is $O(D(1+\frac{1-a^{L}}{1-a}))$, where $L$ is the height of the resulting hierarchical community structure and $a$ is the average ratio of the number of nodes between two consecutive layers, with $0<a<1$.
@@
ArchRAG first computes the embedding of the query, which takes $O(e)$ time. For each layer, querying one nearest neighbor takes no more than $O(\log(n))$ time, similar to the proof in (Malkov and Yashunin 2018). In the Adaptive filtering-based generation, the content of each query is analyzed and inferred, requiring $O(LkI)$ time. Therefore, the total time for the online retrieval query is $O(e+LkI+Lk\log(n))$. ∎
@@
Given a C-HNSW with $L$ layers constructed from a large text corpus or a large set of text documents, the number of tokens used for a single online retrieval query in ArchRAG is $O(kL(c+P)))$, where $k$ is the number of neighbors retrieved at each layer, $c$ is the average token of the retrieved content, and $P$ is the token of the prompt.
@@ Appendix B Experimental details
$\bullet$ Metrics for specific QA tasks. We choose accuracy as the evaluation metric based on whether the gold answers are included in the model’s generations rather than strictly requiring an exact match, following (Schick et al. 2024; Mallen et al. 2022; Asai et al. 2023). This is because LLM outputs are typically uncontrollable, making it difficult for them to match the exact wording of standard answers. Similarly, we choose recall as the metric instead of precision, as it better reflects the accuracy of the generated responses. Additionally, when calculating recall, we adopt the same approach as previous methods (Gutiérrez et al. 2024; Asai et al. 2023): if the golden answer or the generated output contains “yes” or “no”, the recall for that question is set to 0. Therefore, the recall metric is not perfectly correlated with accuracy.
@@
$\bullet$ Metrics for abstract QA tasks. Following existing works, we use an LLM to generate abstract questions, with the prompts shown in Figure 17, defining ground truth for abstract questions, particularly those involving complex high-level semantics, poses significant challenges. We build on existing works (Edge et al. 2024; Guo et al. 2024) to address this and adopt an LLM-based multi-dimensional comparison method (including comprehensiveness, diversity, empowerment, and overall). We employ a robust LLM, specifically GPT-4o, to rank each baseline against our method. Figure 18 shows the evaluation prompt we use.
  1. 1.
    Calinski-Harabasz Index (CHI) (Caliński and Harabasz 1974): A higher value of CHI indicates better clustering results because it means that the data points are more spread out between clusters than they are within clusters. It is an internal evaluation metric where the assessment of the clustering quality is based solely on the dataset and the clustering results and not on external ground-truth labels. The CHI is calculated by between-cluster separation and within-cluster dispersion:
[…]
  2. 2.
    Cosine Similarity (Sim) (Charikar 2002): Cosine similarity is a measure of similarity between two non-zero vectors defined in an inner product space. In this paper, for each cluster, we calculate the similarity between the centroid of this cluster and each node in this cluster:
    $Sim=\frac{1}{N}\sum^{C}_{i=1}\sum_{x\in C_{i}}Cosine(x,c_{i}).$ (4)
    NN is the number of nodes. $C$ is the number of clusters. $C_{i}$ is the $i$-th cluster. $c_{i}$ is the centroid of cluster $C_{i}$. $x$ is the feature of the target node.
    $Cosine(x,y)=\frac{\sum_{i=1}^{n}x_{i}y_{i}}{\sqrt{\sum_{i=1}^{n}x_{i}^{2}}\sqrt{\sum_{i=1}^{n}y_{i}^{2}}}.$ (5)
Implementation details
We implement our ArchRAG in Python, while C-HNSW is implemented in C++ and provides a Python interface for integration. We implement C-HNSW using the FAISS framework and employ the inner product metric to measure the proximity between two vectors. All the experiments were conducted on a Linux operating system running on a machine with an Intel Xeon 2.0 GHz CPU, 1024GB of memory, and 8 NVIDIA GeForce RTX A5000 GPUs, each with 24 GB of memory. All methods utilize 10 concurrent LLM calls, and to maintain consistency, other parallel computations in the method, such as embedding calculations, also use 10 concurrent threads. Figure 19 demonstrates the prompt used in Adaptive filtering-based generation. Please refer to our repository (https://anonymous.4open.science/r/H-CAR-AG-1E0B/) to view the detailed prompts.
$\bullet$ Details of clustering methods. The graph augmentation methods we choose are the KNN algorithm, which computes the similarity between each node, and CODICIL (Ruan, Fuhry, and Parthasarathy 2013), which selects and adds the top similar edges to generate better clustering results. In the KNN method, we set the $K$ value as the average degree of nodes in the KG.
Additional experiments
$\bullet$ Efficiency of hierarchical search (C-HNSW). To evaluate the efficiency of C-HNSW, we conduct experiments on a synthetic hierarchical dataset comprising 11 layers. The bottom layer (Layer 0) contains 10 million nodes, and the number of nodes decreases progressively across higher layers by randomly dividing each layer’s size by 3 or 4. The top layer (Layer 10), for example, contains only 74 nodes. This hierarchical structure simulates the process of LLM-based hierarchical clustering. Each node is assigned a randomly generated 3072-dimensional vector, simulating high-dimensional embeddings such as those produced by text or image encoders (e.g., text-embedding-3-large, used in ChatGPT, can generate 3072-dimensional vectors, and text-embedding-v3 can generate 1024-dimensional vectors). For each layer, we generate 200 random queries and compute the top-5 nearest neighbors for each query. Both C-HNSW and Base-HNSW are configured with identical parameters: $M=32$, $efSearch=100$, and $efConstruction=100$. Importantly, our method maintains comparable retrieval accuracy to Base-HNSW, with recall of 0.5537 and 0.6058, respectively.
[…]
8
012345678910Avg.05510101515Leveltime (s)
Figure 8: C-HNSW and Base-HNSW query efficiency.
We also conducted experiments on a synthetic dataset of 1024-dimensional vectors, keeping all other parameters unchanged. The results are shown in Figure 9. As 1024 dimensions are more representative of commonly used high-dimensional embeddings, our method still achieves over 5× speedup in the best case and an average speedup of 3× compared to the baseline.
9
012345678910Avg.05510101515Leveltime (s)
Figure 9: C-HNSW and Base-HNSW query efficiency on 1024-dimensional vector dataset.
$\bullet$ Efficiency of indexing phrase. Figure LABEL:fig:index-efficiency shows the index construction time and token usage for different methods. The cost of building an index for ArchRAG is similar to that of GraphRAG, but due to the need for community summarization, both take a higher time cost and token usage than HippoRAG.
LABEL:eff_index
Table 5: Comparing ArchRAG with other RAG methods on the specific QA tasks under different LLM backbone models.
LLM backbone Methods Multihop-RAG HotpotQA
(Accuracy) (Recall) (Accuracy) (Recall)
Llama3.1-8B Vanilla RAG 58.6 31.4 50.6 56.1
HippoRAG 38.9 19.1 51.3 56.8
RAPTOR 59.1 34.1 N/A N/A
ArchRAG 68.8 37.2 65.4 69.2
GPT-3.5-turbo Vanilla RAG 65.9 32.8 60.7 65.9
HippoRAG 68.9 31.4 58.0 62.3
RAPTOR 64.4 34.6 N/A N/A
ArchRAG 67.2 31.5 62.8 65.0
GPT-4o-mini Vanilla RAG 71.4 32.8 68.2 70.1
HippoRAG 70.5 31.6 65.0 68.5
RAPTOR 70.1 32.6 N/A N/A
ArchRAG 77.3 33.8 69.9 73.8
To further demonstrate the effectiveness of ArchRAG, we conduct the following experiments:
$\bullet$ Effectiveness of LLM backbones. Given the limited budget, we restrict our evaluation to GPT-4o-mini and GPT-3.5-turbo as the LLM backbones, and compare a representative subset of strong RAG methods on the HotpotQA and Multihop-RAG datasets. As strong LLMs with hundreds of billions of parameters (e.g., GPT-3.5-turbo) possess enhanced capabilities, our proposed ArchRAG may also benefit from performance improvement. As shown in Table 5, the results of Llama 3.1-8B are similar to those of GPT-3.5-turbo, as Llama3.1’s capabilities are comparable to those of GPT-3.5-turbo (Dubey et al. 2024). GPT-4o-mini performs better than other LLM backbones because of its exceptional reasoning capabilities.
Besides, we have compared several strong RAG baselines under different LLM backbones. As LLMs’ parameters and reasoning capabilities increase, all RAG approaches benefit from performance gains, especially HippoRAG. ArchRAG consistently achieves state-of-the-art performance across most settings.
$\bullet$ Community quality of different clustering methods. We evaluate the community quality of our proposed LLM-based hierarchical clustering framework using four clustering algorithms (weighted Leiden, Spectral Clustering (Von Luxburg 2007), SCAN (Xu et al. 2007), and node2vec (Grover and Leskovec 2016) with KMeans), combined with two graph augmentation techniques (the KNN algorithm and CODICIL (Ruan, Fuhry, and Parthasarathy 2013)). The resulting communities are assessed using the Calinski-Harabasz Index (CHI) (Caliński and Harabasz 1974) and Cosine Similarity (Sim) (Charikar 2002), where higher values indicate better quality. Further details on the clustering implementation and evaluation metrics can be found in the appendix. Figures LABEL:fig:ch and LABEL:fig:sim show that combining KNN or CODICIL with the weighted Leiden algorithm significantly enhances community detection quality.
LABEL:quality_ch
LABEL:quality_sim
$\bullet$ Effectiveness of our attributed clustering algorithm. To further demonstrate the effectiveness of our attributed clustering algorithm, we evaluate the quality of communities (i.e., CHI and Cosine Similarity) generated by our attributed clustering algorithm compared to those produced by the Leiden algorithm, which is used in GraphRAG for structural clustering. As shown in Table 7, our attribute-based clustering consistently yields higher-quality communities.
Table 6: Comparing ArchRAG with other RAG methods on the RAG-QA Area dataset. Each entry denotes the win ratio and win + tie ratio of the corresponding method against the ground-truth annotations, based on LLM evaluation.
Method Lifestyle Recreation Science Technology Writing
Vanilla RAG 17.5 / 20.5 17.0 / 25.0 32.5 / 37.0 28.5 / 34.0 15.0 / 16.5
HippoRAG 26.5 / 26.5 29.5 / 30.5 49.5 / 49.5 42.0 / 42.5 21.0 / 21.5
RAPTOR 17.0 / 19.5 18.5 / 24.5 39.0 / 45.0 33.0 / 35.5 25.0 / 26.5
ArchRAG 49.5 / 50.0 41.5 / 41.5 56.0 / 56.0 59.0 / 59.5 45.0 / 45.0
Table 7: Comparison of Community Quality between Our Attributed Clustering Method and Leiden
Method Multihop-RAG HotpotQA
(CHI) (Sim) (CHI) (Sim)
Leiden 3.02 0.71 3.42 0.71
Ours 4.68 0.89 4.82 0.88
$\bullet$ More experiments on the additional dataset. We further conduct experiments on the RAG-QA Arena dataset (Han et al. 2024), a high-quality, multi-domain benchmark featuring human-annotated, coherent long-form answers. To the best of our capability, we use publicly available data from five domains (including lifestyle, recreation, science, technology, and writing), selecting 200 questions per domain. Following prior work, we employ LLMs as evaluators to compare the RAG-generated responses with ground-truth answers in terms of win ratio and win + tie ratio. We evaluate the top-performing methods (including Vanilla RAG, HippoRAG, RAPTOR, and our ArchRAG) and present the results in terms of the win ratio and win + tie ratio against the ground-truth annotations in the table below. As shown in Table 6, ArchRAG consistently achieves state-of-the-art performance across all evaluated settings.
LABEL:topk_leg
$\bullet$ Case study. We present an additional case study from Multihop-RAG. As shown in Figure 15, only our method generates the correct answer, while others either provide incorrect or irrelevant information. We only show the core output for brevity, with the remaining marked as “¡/¿”. We also show the retrieval and adaptive filtering process of ArchRAG in Figure 16. The results demonstrate that ArchRAG effectively retrieves relevant information and filters out noise, leading to a correct final answer.
Table 8: Distribution of HippoRAG’s ER Errors
Datasets Null Entity Rate
Low-Quality
Entity Rate
Multihop-RAG 1.3% 11.9%
HotpotQA 5.0% 15.8%
$\bullet$ Discussion of the performance of other graph-based RAG methods. On the Multihop-RAG dataset, HippoRAG performs worse than retrieval-only methods while outperforming retrieval-based methods on the HotpotQA dataset. This is mainly because, on the HotpotQA dataset, passages are segmented by the expert annotators; that is, passages can provide more concise information, whereas on Multihop-RAG, passages are segmented based on chunk size, which may cause the LLM to lose context and produce incorrect answers. Besides, HippoRAG also suffers from inaccurate entity recognition. In addition, we also provide a brief analysis of why some graph-based RAG methods underperform. Specifically, they often fail to retrieve relevant information during the search process. For example:
  • GraphRAG: The suboptimal performance of GraphRAG has been widely observed in research. For instance, Zhang et al. (Zhang et al. 2024a) find that GraphRAG tends to respond with ”I don’t know” when retrieving irrelevant content, indicating that it prefers not to give a concrete answer. PIKE-RAG (Wang et al. 2025) identifies that GraphRAG tends to echo the query and include meta-information about answers within its graph structure. Through manual verification, we confirm the presence of similar issues in our experimental results. In particular, GraphRAG-Local Search also frequently returns irrelevant content.
  • HippoRAG/LightRAG: We attribute part of HippoRAG’s suboptimal performance to entity recognition issues. HippoRAG first extracts relevant entities from questions and then retrieves chunks based on these entities. However, it sometimes fails to identify appropriate entities (Entity Recognition Errors, ER Errors), and we conduct corresponding experiments, as shown in Table 8. For example, on HotpotQA, HippoRAG fails to detect entities in 5.0% of questions and identifies low-quality ones in 15.8% (e.g., less than 3 characters, numeric-only terms, or sentence entities with more than 5 words). These issues make it difficult to match accurate entities in the KG, resulting in irrelevant retrieval and suboptimal performance. Similar issues also impair LightRAG’s performance because it requires accurate topic identification from questions. We also observe that as the LLM’s capability improves, the performance of these methods increases significantly, as shown in the experiment Effectiveness of LLM backbone.
Question & Ground-truth Question: Who is the individual associated with generative AI technology that was reportedly ousted from a leading AI organization, but is recognized for brilliance and generosity in the industry, and is currently planning to launch a new venture according to reports from TechCrunch and Fortune? Ground-truth: Sam Altman
Responses of Different Methods Zero-shot Jesse Schell.
He is the CTO of Calm and was the lead designer on the “Star Wars” theme park ride at Disney.
  Vanilla RAG It seems like there are multiple articles and topics discussed in this text. Here’s a brief summary of each section:
1.Robotics predictions. The article discusses the future of robotics, with predictions made by Brian Heater. $\cdots$
2.TechCrunch’s Week in Review. This section summarizes various tech industry happenings. $\cdots$
3.Bumble gets a new CEO. Bumble announces a change in leadership. $\cdots$
  HippoRAG There is no question in the provided text. It appears to be a collection of news articles and updates related to OpenAI’s ChatGPT. $\cdots$
  GraphRAG-Global Key Points and Implications
The individual associated with generative AI technology who was reportedly ousted is Andrew Ng. $\cdots$
According to multiple analysts, Andrew Ng [Data: Reports (5, 6)] $\cdots$
Implications of the New Venture $\cdots$
  ArchRAG Sam Altman.
This information is supported by multiple analysts’ reports, which highlight Sam Altman’s involvement in various AI-related projects. The direct answer is simply “Sam Altman”. $\cdots$
Note: I have not included any information from Analyst 2, 3, or 7 as their reports are not relevant. $\cdots$
Figure 15: Case study of responses by different RAG methods on a question from the Multihop-RAG dataset. RED TEXT indicates incorrect content, CYAN TEXT denotes correct content, and $\cdots$ marks omitted irrelevant parts.
Top Retrieved Entities:
- Entity 1058: OPENAI
Description: OpenAI is a tech company founded by Sam Altman ...
- Community 4097: Microsoft, OpenAI, and AI Regulation
Summary: Discussion between Satya Nadella and Sam Altman ...
Adaptive Filtering Result:
- Sam Altman is among the backers of an AI startup.
- Score: 80.0 (Reports: 1) ...
Figure 16: ArchRAG Retrieval & Filtering Output. ... marks omitted irrelevant parts.
@@
Prompt for generating abstract questions Prompt: Given the following description of a dataset: {description} Please identify 5 potential users who would engage with this dataset. For each user, list 5 tasks they would perform with this dataset. Then, for each (user, task) combination, generate 5 questions that require a high-level understanding of the entire dataset. Output the results in the following structure: - User 1: [user description] - Task 1: [task description] - Question 1: - Question 2: - Question 3: - Question 4: - Question 5: - Task 2: [task description] - Task 5: [task description] - User 2: [user description] - User 5: [user description] Note that there are 5 users and 5 tasks for each user, resulting in 25 tasks in total. Each task should have 5 questions, resulting in 125 questions in total. The Output should present the whole tasks and questions for each user. Output:
Figure 17: The prompt for generating abstract questions.
@@
        "Explanation": "[Briefly summarize why this answer is the overall
        winner]"
@@
Figure 18: The prompt for the evaluation of abstract QA.
Prompt for Adaptive filtering-based generation Filter Prompt: # Role You are a helpful assistant responding to questions about data in the tables provided. # Goal Generate a response consisting of a list of key points that respond to the user’s question, summarizing all relevant information in the input data tables. You should use the data provided in the data tables below as the primary context for generating the response. If you don’t know the answer or if the input data tables do not contain sufficient information to provide an answer, just say so. Do not make anything up. Each key point in the response should have the following element: Description: A comprehensive description of the point. Importance Score: An integer score between 0-100 that indicates how important the point is in answering the user’s question. An ‘I don’t know’ type of response should have a score of 0. The response should be JSON formatted as follows: